OpenStudy (anonymous):

Does the equation x^2+7y^2=1 define y as a function of x?

OpenStudy (savvy):

\[y = 1/7(\sqrt{1-x ^{2}})\]

OpenStudy (anonymous):

Thank you!!! Your awsome! I was not sure if I was soing the problem right or not! =)

OpenStudy (anonymous):

Solving for y we get:\[y = \pm \sqrt{\frac{1-x^2}{7}}\]If we take both values for the square root, it's not a function, because of the vertical line test. If we choose either the positive square root or the negative square root, we get a function. As the question does not state either, it's not.

OpenStudy (anonymous):

thank you, execlent work! That was my first question and Im very impressed at how fast that was!

OpenStudy (anonymous):

The visual graphs are great! Thank you sooo much! I am having such a hard time in my collage algebra right now. I've done great in everything but this!